A person's level of blood glucose and diabetes are closely related. Let x be a random variable measured in milligrams of glucose per deciliter (1/10 of a liter) of blood. Suppose that after a 12-hour fast, the random variable x will have a distribution that is approximately normal with mean μ = 87 and standard deviation of σ = 30. What is the probability that, for an adult after a 12-hour fast, x is between 96 and 138? Round your answer to the nearest thousandth.

Group of answer choices

0.169

0.787

0.338

0.955

0.045
Answer #1

z score for 96 = (96 - 87)/30 = 0.30

z score for 138 = (138 - 87)/30 = 1.7

Hence,

P(X is between 96 and 138)

= P(0.3 < z < 1.7)

= 0.338